OpenStudy (anonymous):

An automobile traveling 10 mi/hr has wheels 36 inches in diameter. If the car is brought to a stop uniformly in 74 turns of the wheels, what is the magnitude of the angular acceleration of the wheels (in radians/second^2)? So i've found just the radians/sec which is 9.77. After that i'm not quite sure where to go. I'm I supposed to use those "kinematic" equations?

OpenStudy (shamim):

Can u tell me retardation a=?

OpenStudy (shamim):

Use v^2=u^2-2as a=?

OpenStudy (shamim):

Given v=0 u=10mi/h=?inch/sec s=74turn*circumference of wheel=74*pi*diameter of wheel=74*pi*36=?inches

OpenStudy (shamim):

U know a=alpha*r Alpha=?

OpenStudy (shamim):

Alpha=angular acceleration

OpenStudy (shamim):

r=radius of wheel=36/2=?inches